(Computing income taxes )

The William B. Waugh Corporation is a regional Toyota dealer. The firm sells new and used trucks and is actively involved in the parts business. During the most recent year, the company generated sales of $ 3.05 million. The combined cost of goods sold and the operating expenses were $2.02 million. Also, $387000 in interest expense was paid during the year. Calculate the corporation's tax liability by using the corporate tax rate structure in the popup window,

Corporate Tax Rates

​ 15%

​$0long dash—​$50,000

​ 25%

​$50,001long dash—​$75,000

​ 34%

​$75,001long dash—​$10,000,000

​ 35%

over​ $10,000,000

  Additional​ surtax:

  

bullet•

​5% on income between​ $100,000 and​ $335,000
  

bullet•

​3% on income between​ $15,000,000 and​ $18,333,333
Answer #1

taxable income= Sales-cost of goods sold-operating expenses-interest expenses
=(3.05*10^6)-(2.02*10^6)-387000
=643000
tax liability based on table provided below:
=(50000*15%)+((75000-50000)*25%)+((643000-75000)*34%)
=206870
5% additional=((335000-100000)*5%)=11750
Total corporate tax liability=206870+11750=218620